2. Suppose you were testing US The sample is large (n 61) and the variance is known. (a) Find the critical value(s) corresponding to α-0.06 (b) You find that z-_1.6. Based on your critical value, what decision do you make regarding the null hypothesis (i.e. do you Reject Ho or Do Not Reject Ho)?
Answer #1

a) This is a two tailed test with known population variance and large sample size. Hence,

Critical values for 0.06 level of significance = -1.88, 1.88

b) This is a two tailed test. Hence,

p - Value = 2*P(z < -1.6) = 2*0.0548 = 0.1096

SInce this p - value is greater than 0.06, we do not reject Ho